Keller Company makes two models of battery-operated boats, the Sandy Beach and the Rocky River. Basic production information follows:

Sandy Beach Rocky River
Direct materials cost per unit $ 20.00 $ 27.70
Direct labor cost per unit 14.90 17.40
Sales price per unit 83.10 105.00
Expected production per month 1,200 units 980 units

      

Keller has monthly overhead of $11,245, which is divided into the following cost pools:

Setup costs $ 1,920
Quality control 6,077
Maintenance 3,248
Total $ 11,245

The company has also compiled the following information about the chosen cost drivers:

Sandy Beach Rocky River Total
Number of setups 15 15 30
Number of inspections 110 405 515
Number of machine hours 1,450 1,450 2,900

Required:
1.
Suppose Keller uses a traditional costing system with machine hours as the cost driver. Determine the amount of overhead assigned to each product line. (Do not round intermediate calculations and round your final answers to the nearest whole dollar amount.)




2. Calculate the production cost per unit for each of Keller’s products under a traditional costing system. (Round your intermediate calculations and final answers to 2 decimal places.)



3. Calculate Keller’s gross margin per unit for each product under the traditional costing system. (Round your intermediate calculations and final answers to 2 decimal places.)



4. Select the appropriate cost driver for each cost pool and calculate the activity rates if Keller wanted to implement an ABC system.(Round your answers to 2 decimal places.)
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Hickory
Workings for Answer 1
Calculation of Plant wide OH Rate Amount $
Manufacturing overhead         11,245.00 A
Machine Hours            2,900.00 B
Plant wide OH Rate                   3.88 C=A/B
Answer 1 Sandy Rocky Note
Machine Hours            1,450.00       1,450.00 D
Predetermined OH Rate                   3.88               3.88 See C
Overhead assigned           5,623.00       5,622.00 E=C*D
Workings for Answer 2 Sandy Rocky
Overhead assigned            5,623.00       5,622.00 See E
Expected production per month            1,200.00           980.00 F
Overhead assigned per unit                   4.69               5.74 G=E/F
Answer 2 Sandy Rocky
Direct Materials                 20.00             27.70
Direct Labor                 14.90             17.40
Overhead assigned per unit                   4.69               5.74 See G
Production cost per unit                 39.59             50.84
Answer 3 Sandy Rocky
Sell price per unit                 83.10           105.00
Production cost per unit                 39.59             50.84
Gross margin per unit                 43.51             54.16
Answer 4
Calculation of ABC rates for overhead H I J=H/I
Activity cost pool Cost Cost Driver Activity usage ABC rates
Setup costs            1,920.00 Number of setups                 30.00             64.00
Quality control            6,077.00 Number of inspections               515.00             11.80
Maintenance            3,248.00 Number of Machine-hours               2,900.00                1.12
        11,245.00